OpenStudy (anonymous):

rationalize the numerator... sqrt of 5/7

OpenStudy (anonymous):

\[\frac{\sqrt{5}}{7}=\frac{\sqrt{5}}{7}\times \frac{\sqrt{5}}{\sqrt{5}}\] \[=\frac{5}{7\sqrt{5}}\]

OpenStudy (anonymous):

that isn't a choice for an answer :[

OpenStudy (anonymous):

a. 5/sqrt7 b. 7/sqrt7 c. 5/sqrt35 d. sqrt35/7